# I. Márquez 11    C. Boisson 22    M. Joly 22    D. Pelat 22    F. Durret 33

### v1

@@ Abstract
Context. The relationship between an active galactic nucleus (AGN) and its host galaxy is still far from being understood. Properties of the host galaxies of Seyfert nuclei, such as luminosity concentration, morphological type, metallicity, and age of the stellar population are expected to be related with nuclear activity – either at the epoch of galaxy formation or in the present days via feeding of the central black-hole.
Aims. In this paper we investigate whether stellar ages and metallicities are linked to the activity within the nucleus in a sample of AGN of various types.
Methods. Our sample includes seven AGN, from Seyfert 1 to LINERs, observed with VLT/ISAAC and VLT/SINFONI. Based on an inverse method using a stellar library, we analyse H band infrared spectra, in a wavelength region devoid of emission lines, at a spectral resolution of $\approx$3000, in the central few 100 pc. HST images are used to visualise the regions defined in each galaxy.
Results. For each galaxy, we give the results of the spectral synthesis, in particular the percentages of the stellar, power law and blackbody continua, and the percentages of various stellar types that account for the stellar lines.
Conclusions. Out of the seven galaxies, three show strong and recent star formation in the inner 100 pc, while no star formation is detected in the three genuine Seyfert 2 galaxies. Beyond a radius of 100 pc, all show more or less recent star formation. Moreover we can conclude that the star formation history of the inner nucleus is highly heterogeneous.
@@ 1 Introduction
An important yet unsolved issue in understanding the Seyfert phenomenon is the relation between the host galaxy and the properties of activity within the nucleus. It is now accepted that nuclear star formation and central engine activity can coexist in the inner regions of the active galactic nuclei (AGN) host galaxy, as derived from the strong relation linking the mass of super-massive black holes with the velocity dispersion of their host galaxy bulge (e.g., Gebhardt et al., 2000; Kormendy & Ho, 2013). The accelerated star formation caused by dynamic instabilities that trigger and/or fuel nuclear activity could result in an overabundance of giant, supergiant, and supermetal-rich (SMR) stars (e.g., Scoville et al., 2023, and references therein).
Stellar population syntheses of the cores of AGN provide information on stellar masses, mean stellar ages, abundances, dust content, and insight on their star-formation history. Such informations are derived from spatially unresolved galaxy spectra and thus reflect global averages over properties that are known to vary across the galaxies, but still provide valuable insights. Observations of the cores of nearby active galaxies allow to exhibit the complex star formation history in active galaxies (e.g., Terlevich et al., 1990; Storchi-Bergmann et al., 2001; Diniz et al., 2017; Mallmann et al., 2018; Dametto et al., 2019). Studies in the optical and infrared domains have indicated that young or intermediate age populations can contribute significantly to AGN central spectral emission (e.g., Oliva et al., 1999; Boisson et al., 2000; Chapman et al., 2000; Boisson et al., 2004; García-Bernete et al., 2015; Thomas et al., 2018; Burtscher et al., 2021; Dahmer-Hahn et al., 2022; Riffel et al., 2024, and references therein).
To study the properties of the stellar population in the inner few tens of parsecs, all studies face the problem of disentangling the stellar population from the other components: the continuum emission by the AGN, the emission from heated dust, and galactic extinction. Contrary to the optical spectral range, near-IR allows access to regions highly obscured by dust, and exhibits many stellar features that allow to constrain the average spectral type, luminosity class, and metallicity markers characteristic of a large variety of stellar populations (e.g., Origlia et al., 1993; Dallier et al., 1996; Ivanov et al., 2004; Stephens & Frogel, 2004).
Since part of the near-IR H-band is free of the strongest emission lines due to the AGN, it is a well suited domain for the determination of the stellar content of central regions of Seyfert galaxies, including Seyfert 1s, as long as the spectral resolution is high enough (e.g., Dallier et al., 1996; Frémaux et al., 2006; Cesetti et al., 2009).
@@
In Section 2 of the paper we briefly present the sample of seven nearby AGN observed with the VLT and summarise the data reduction. The spectral population synthesis code and its implementation are presented in Section 3. Results for individual AGN are presented in Section 4 and discussed in a global context in Section 5.
@@ 2 The data
To study the differential stellar population in the core of AGN we assembled a sample of seven sources typical of their class of activity, three Seyfert 1 galaxies (MCG-6-30-15, NGC 2992 and NGC 3783) and four Seyfert 2s (NGC 2110, NGC 3185, NGC 6221 and NGC 7582). Five galaxies observed with the VLT/ISAAC spectrograph, were previously presented in Boisson et al. (2002) and are complemented by two more galaxies (this paper) observed with the same equipment. We combined these data with VLT/SINFONI spectra kindly provided by Sandra Raimundo for MCG-6-30-15 (Raimundo et al., 2013).
HST archival images were used to visualise the different regions as selected for each galaxy. Informations on the seven galaxies of this study are given in Table 1.
In this paper we assume H0 = 70 km s-1 Mpc-1, $\Omega_{m}$=0.3, $\Omega_{\Lambda}$=0.7.
@@
The columns are: 1) name, 2) Seyfert type, 3) morphological type, 4) and 5) coordinates, 6) redshift, 7) Galactic E(B-V), 8) slit position angle, 9) HST filter, 10) distance, 11) scale in pc per arcsecond. Galactic E(B-V) values are from Schlafly & Finkbeiner (2011) as given by the Dust Extinction tool provided by IRSA (https://irsa.ipac.caltech.edu/applications/DUST/).
@@ 2.1 VLT/ISAAC data
The slits were always centred on the infrared (IR) galaxy nucleus. Their position angles are provided in Table 1.
The data reduction is fully described in Boisson et al. (2002), including correction for electrical ghosts, flat-field, distortions along and perpendicular to the slit, and wavelength calibration. Early type stellar templates were used for the correction of telluric absorption. The spectra were extracted by using logarithmic flux profiles to define a nucleus and regions around it. All of the spectra displayed are shifted to the rest wavelength.
Whenever spectra on each side of the nucleus are very similar, a pseudo ring spectrum is created by summing both individual spectra, allowing a better signal to noise ratio.
@@ 2.3 HST images
Archival WFPC2 HST images are available for all the galaxies (see the HST filters in Table 1). The reddest high quality images were used. In order to enhance morphological features in the innermost few arcseconds of the galaxies, we created sharp-divided images by dividing the original images by their smoothed counterparts (Márquez et al., 2003). This enhances any structural inhomogeneities that could affect the stellar populations, such as a dust component.
The spectral regions extracted from ISAAC and SINFONI observations are drawn on the images, the slit width is 1″. North is up and East to the left.
@@ 3 Stellar population synthesis
The analysis of the stellar populations inside unresolved cores of active galaxies has been the aim of several studies for some time now. All studies face the problem of disentangling the stellar population from all the other components: the continuum emission by the AGN, the emission from the heated dust and galactic extinction.
@@
The near-infrared range is a very useful domain to study the stellar content of active galaxies. In particular the H-band is well suited as the non-stellar contribution (mainly reprocessed nuclear emission by dust) is much less than in the K-band window. The H-band is also the region in which emission from cool stars peaks, and the spectrum is rich in luminosity sensitive stellar absorption lines, leading to the age of the stellar population (e.g., Dallier et al., 1996; Frémaux et al., 2006). Extinction by dust is lower there than for the shorter wavelength J-band.
@@
Since part of the infrared H-band is free of the strongest AGN emission lines, it may allow for the determination of stellar populations in the very internal regions of Seyfert galaxies in presence of a diluting continuum, as is the case for Seyfert 1s. In particular the spectral range considered in this paper, covering 1.57 to 1.64 $\mu$m, includes many metallic stellar features (FeI, MgI, SiI, CaI, NiI, CO and OH). The spectral variations of the stellar component are generally complex, the age and the metallicity of the stars may have similar influences on the integrated spectrum. Therefore, it is crucial to use as many observables as possible to decrease the degeneracy. Indeed the near-IR CaII triplet and the IR CO and Si indices can be strong not only in supergiant stars but also in metal-rich giants; colour gradients can be caused either by changing stellar populations or by dust effects.
@@
There are essentially two different approaches to differential population synthesis. The “direct” approach, also called “evolutive synthesis method” (e.g., Tinsley, 1972; Bruzual & Charlot, 2003; Le Borgne et al., 2004; Vázquez & Leitherer, 2005), compares the spectrum at each time step of an evolving stellar population with the spectrum of the studied galaxy. The other methods aim to extract the stellar components from the spectrum of the astrophysical object through a mathematical inversion of the information content of the observed spectrum. These so-called “inverse” methods use building blocks of either stellar spectra (e.g., Faber, 1972; Pickles, 1985; Pelat, 1997, 1998; Moultaka & Pelat, 2000; Frémaux et al., 2007) or simple stellar populations, SSP (e.g., Bica, 1988; Vergely et al., 2002; Cid Fernandes et al., 2005; Gomes & Papaderos, 2017; Wilkinson et al., 2017; Cid Fernandes, 2018). In this paper, we use a similar method based on the minimization of a quadratic form subject to constraints in the form of inequalities and equalities, developed and thoroughly tested by Frémaux et al. (2007).
@@
Taking the flux received in each spectral pixel as the observable quantity to fit increases the number of observables, allowing a greater sampling for the stellar composition of the studied galaxies, and alleviating the difficulties to determine the level and shape of the continuum. The specificity of the method is to take the non stellar parameters (reddening, dust emission and AGN non-stellar continuum) directly into account in the synthetic distance to be minimised. These non-stellar parameters introduce non-linearity into the objective function, expressed as a quadratic distance to minimise. The lower this value, the closer from the observation the synthetic spectrum.
@@
Due to degeneracy of these components, mainly in low S/N ($<$ 40), the non-stellar parameters are bound to values commonly deduced from astrophysical considerations. Dust grains with a temperature below 200 K mainly contribute to the flux spectrum farther in the infrared, so their contribution in our wavelength range is too small to be detected. A temperature of 1500 K is the maximum possible for dust grains as it corresponds to the sublimation temperature of the most refractory dust material. Most of the active galaxies have an intrinsic reddening E(B-V) smaller than 3. For each galaxy the minimum allowed galactic extinction was fixed to the Galactic value on the line of sight (Schlafly & Finkbeiner, 2011). The AGN continuum power law index $\alpha$ is the most difficult non stellar parameter to estimate. Indeed, in our very small wavelength domain, variations of $\alpha$ only induce a very weak variation in the slope of the fitted continuum, as explained in Frémaux et al. (2007). We therefore fixed $\alpha$ to the typical value of 1.5 for AGN111the power law is defined by $\nu^{-\alpha}$ (Vanden Berk et al., 2001). However, these bounds can easily be modified for the synthesis of individual objects.
@@
Such a method is inherently prone to errors due to the presence of noise in the data and potential degeneracies inherent to the studied problem. To assess the impact of these errors on the estimated parameters, we performed a series of Monte Carlo simulations. Gaussian random noise with zero mean and a standard deviation of 10% of the observed galactic spectrum value was added to each pixel of the spectrum. This corresponds approximately to the signal-to-noise ratio of the data. For each object, 300 independent Monte Carlo trials were conducted. To mitigate the influence of outliers, 38% of the highest and 38% of the lowest values for each parameter were excluded. This percentage corresponds to the optimal truncated estimator for the worst-case scenario, i.e. noise with undefined mean (Cauchy noise). The standard deviation of the remaining values provides the error bars for the parameters obtained via the minimization method.
@@
Whatever the approach in stellar syntheses, a spectral database is needed to build a synthetic galactic spectrum. The stellar libraries come in two flavours, either built with theoretical or empirical spectra. Though theoretical libraries could allow a wide range and combination of stellar parameters, it requires exhaustive model atmospheres that depend on atomic and molecular line opacity data bases not always available. Though theoretical libraries could allow a wide range and combination of stellar parameters, it requires exhaustive model atmospheres that depend on atomic and molecular line opacity data bases not always available.
@@
As demonstrated by Frémaux et al. (2006), the comparison between observed and computed spectra deteriorates for decreasing temperature, where several molecular as well as metallic absorption lines are still missing in atmospheric models. Working in the infrared H-band, we thus use an empirical stellar library. It should be noted that each individual spectral type in a library encompasses a significantly large domain of temperature, gravity, and metallicity, because one cannot have as many types of stars in the database as those forming a galaxy.
@@
In the text, we define as young, intermediate and old stellar populations those with stars with ages $\leq$ 107 yrs (supergiant stars), around 108 yrs (A0-F2V), and older than 109 yrs (dwarfs and giants) respectively. The supergiant stars are indeed used as proxy for a young population since no stars hotter than A type can be considered; the contribution of O and B type stars is negligible in the infrared, compared to cold or intermediate stars (A, F spectral types).
For more details on the spectral fitting code and database see Frémaux et al. (2007).
Table 2: Stellar parameters

### v2

@@ Abstract
Context. The relationship between an active galactic nucleus (AGN) and its host galaxy is still far from being understood. Properties of the host galaxies of Seyfert nuclei, such as luminosity concentration, morphological type, metallicity, and age of the stellar population, are expected to be related with nuclear activity – either at the epoch of galaxy formation or at the present day via feeding of the central black hole.
Aims. In this paper, we investigate whether stellar ages and metallicities are linked to the activity within the nucleus in a sample of AGNs of various types.
Methods. Our sample includes seven AGNs, from Seyfert 1 to LINERs, observed with VLT/ISAAC and VLT/SINFONI. Based on an inverse method using a stellar library, we analysed H band infrared spectra, in a wavelength region devoid of emission lines, at a spectral resolution of $\approx$3000, in the few central 100 pc. Hubble Space Telescope images were used to visualise the regions defined in each galaxy.
Results. For each galaxy, we give the results of the spectral synthesis, in particular the percentages of the stellar, power law, and black-body continua, and the percentages of various stellar types that account for the stellar lines.
Conclusions. Out of the seven galaxies, three show strong and recent star formation in the inner 100 pc, while no star formation is detected in the three genuine Seyfert 2 galaxies. Beyond a radius of 100 pc, all show more or less recent star formation. Moreover, we can conclude that the star formation history of the inner nucleus is highly heterogeneous.
@@ 1 Introduction
An important yet unsolved issue for understanding the Seyfert phenomenon is the relation between the host galaxy and the properties of activity within the nucleus. It is now accepted that nuclear star formation and central engine activity can coexist in the inner regions of the active galactic nuclei (AGNs) host galaxy, as derived from the strong relation linking the mass of super-massive black holes with the velocity dispersion of their host galaxy bulge (e.g. Gebhardt et al., 2000; Kormendy & Ho, 2013). The accelerated star formation caused by dynamic instabilities that trigger and/or fuel nuclear activity could result in an overabundance of giant, supergiant, and supermetal-rich (SMR) stars (e.g. Scoville et al., 2023, and references therein).
Stellar population syntheses of the cores of AGNs provide information on stellar masses, mean stellar ages, abundances, and dust content, as well as insight on their star-formation history. Such information is derived from spatially unresolved galaxy spectra and thus reflect global averages over properties that are known to vary across the galaxies, but still provide valuable insights. Observations of the cores of nearby active galaxies allow the complex star formation history in active galaxies to be exhibited (e.g. Terlevich et al., 1990; Storchi-Bergmann et al., 2001; Diniz et al., 2017; Mallmann et al., 2018; Dametto et al., 2019). Studies in the optical and infrared (IR) domains have indicated that young or intermediate age populations can contribute significantly to AGN central spectral emission (e.g. Oliva et al., 1999; Boisson et al., 2000; Chapman et al., 2000; Boisson et al., 2004; García-Bernete et al., 2015; Thomas et al., 2018; Burtscher et al., 2021; Dahmer-Hahn et al., 2022; Riffel et al., 2024, and references therein).
To study the properties of the stellar population in the inner few tens of parsecs, all studies face the problem of disentangling the stellar population from the other components: the continuum emission by the AGN, the emission from heated dust, and galactic extinction. Contrary to the optical spectral range, near-IR allows one to access to regions highly obscured by dust, and exhibits many stellar features that permit one to constrain the average spectral type, luminosity class, and metallicity markers characteristic of a large variety of stellar populations (e.g. Origlia et al., 1993; Dallier et al., 1996; Ivanov et al., 2004; Stephens & Frogel, 2004). Since part of the near-IR H-band is free of the strongest emission lines due to the AGN, it is a well-suited domain for the determination of the stellar content of central regions of Seyfert galaxies, including Seyfert 1s, as long as the spectral resolution is high enough (e.g. Dallier et al., 1996; Frémaux et al., 2006; Cesetti et al., 2009).
@@
In Section 2 of the paper, we briefly present the sample of seven nearby AGNs observed with the VLT and summarise the data reduction. The spectral population synthesis code and its implementation are presented in Section 3. Results for individual AGN are presented in Section 4 and discussed in a global context in Section 5.
@@ 2 The data
To study the differential stellar population in the cores of AGNs, we assembled a sample of seven sources typical of their class of activity, three Seyfert 1 galaxies (MCG-6-30-15, NGC 2992 and NGC 3783) and four Seyfert 2s (NGC 2110, NGC 3185, NGC 6221 and NGC 7582). Five galaxies observed with the VLT/ISAAC spectrograph, were previously presented in Boisson et al. (2002) and are complemented by two more galaxies (this paper) observed with the same equipment. We combined these data with VLT/SINFONI spectra kindly provided by Sandra Raimundo for MCG-6-30-15 (Raimundo et al., 2013).
Hubble Space Telescope (HST) archival images were used to visualise the different regions as selected for each galaxy. Informations on the seven galaxies of this study are given in Table 1. In this paper we assume H0 = 70 km s-1 Mpc-1, $\Omega_{m}$=0.3, $\Omega_{\Lambda}$=0.7.
@@
The columns are: 1) name, 2) Seyfert type, 3) morphological type, 4) and 5) coordinates, 6) redshift, 7) Galactic E(B-V), 8) slit position angle, 9) HST filter, 10) distance, 11) scale in parsecs per arcsecond. Galactic E(B-V) values are from Schlafly & Finkbeiner (2011) as given by the Dust Extinction tool provided by IRSA (https://irsa.ipac.caltech.edu/applications/DUST/).
@@ 2.1 VLT/ISAAC data
The slits were always centred on the IR galaxy nucleus. Their position angles are provided in Table 1.
The data reduction is fully described in Boisson et al. (2002), including correction for electrical ghosts, flat-field, distortions along and perpendicular to the slit, and wavelength calibration. Early type stellar templates were used for the correction of telluric absorption. The spectra were extracted by using logarithmic flux profiles to define a nucleus and regions around it. All of the spectra displayed are shifted to the rest wavelength. Whenever spectra on each side of the nucleus were very similar, a pseudo ring spectrum was created by summing both individual spectra, allowing for a better signal to noise ratio.
@@ 2.3 HST images
Archival WFPC2 HST images are available for all the galaxies (see the HST filters in Table 1). The reddest high quality images were used. In order to enhance morphological features in the innermost few arcseconds of the galaxies, we created sharp-divided images by dividing the original images by their smoothed counterparts (Márquez et al., 2003). This enhanced any structural inhomogeneities that could affect the stellar populations, such as a dust component.
The spectral regions extracted from ISAAC and SINFONI observations were drawn on the images, and the slit width is 1″. North is up and east to the left.
@@ 3 Stellar population synthesis
The analysis of the stellar populations inside unresolved cores of active galaxies has been the aim of several studies for some time now. All studies face the problem of disentangling the stellar population from all the other components: the continuum emission by the AGN, the emission from the heated dust, and galactic extinction.
@@
The near-IR range is a very useful domain to study the stellar content of active galaxies. In particular, the H-band is well suited, as the non-stellar contribution (mainly reprocessed nuclear emission by dust) is much less than in the K-band window. The H-band is also the region in which emission from cool stars peaks, and the spectrum is rich in luminosity sensitive stellar absorption lines, leading to the age of the stellar population (e.g. Dallier et al., 1996; Frémaux et al., 2006). Extinction by dust is lower there than for the shorter wavelength J-band.
@@
Since part of the IR H-band is free of the strongest AGN emission lines, it may allow for the determination of stellar populations in the very internal regions of Seyfert galaxies in the presence of a diluting continuum, as is the case for Seyfert 1s. In particular the spectral range considered in this paper, covering 1.57 to 1.64 $\mu$m, includes many metallic stellar features (FeI, MgI, SiI, CaI, NiI, CO, and OH). The spectral variations of the stellar component are generally complex, and the age and the metallicity of the stars may have similar influences on the integrated spectrum. Therefore, it is crucial to use as many observables as possible to decrease the degeneracy. Indeed the near-IR CaII triplet and the IR CO and Si indices can be strong not only in supergiant stars but also in metal-rich giants; colour gradients can be caused either by changing stellar populations or by dust effects.
@@
There are essentially two different approaches to differential population synthesis. The ’direct’ approach, also called the ’evolutive synthesis method’ (e.g. Tinsley, 1972; Bruzual & Charlot, 2003; Le Borgne et al., 2004; Vázquez & Leitherer, 2005), compares the spectrum at each time step of an evolving stellar population with the spectrum of the studied galaxy. The other method aims to extract the stellar components from the spectrum of the astrophysical object through a mathematical inversion of the information content of the observed spectrum. This so-called inverse method uses building blocks of either stellar spectra (e.g. Faber, 1972; Pickles, 1985; Pelat, 1997, 1998; Moultaka & Pelat, 2000; Frémaux et al., 2007) or simple stellar populations (SSPs; e.g. Bica, 1988; Vergely et al., 2002; Cid Fernandes et al., 2005; Gomes & Papaderos, 2017; Wilkinson et al., 2017; Cid Fernandes, 2018). In this paper, we use an inverse method, based on the minimisation of a quadratic form subject to constraints in the form of inequalities and equalities, developed and thoroughly tested by Frémaux et al. (2007).
@@
Taking the flux received in each spectral pixel as the observable quantity to fit increases the number of observables, allowing for a greater sampling for the stellar composition of the studied galaxies, and alleviating the difficulties to determine the level and shape of the continuum. The specificity of the method is to take the non stellar parameters (reddening, dust emission, and AGN non-stellar continuum) directly into account in the synthetic distance to be minimised. These non-stellar parameters introduce non-linearity into the objective function, expressed as a quadratic distance to minimise. The lower this value, the closer from the observation the synthetic spectrum.
@@
Due to degeneracy of these components, mainly in low S/N ($<$ 40), the non-stellar parameters are bound to values commonly deduced from astrophysical considerations. Dust grains with a temperature below 200 K mainly contribute to the flux spectrum farther in the IR, so their contribution in our wavelength range is too small to be detected. A temperature of 1500 K is the maximum possible for dust grains as it corresponds to the sublimation temperature of the most refractory dust material. Most of the active galaxies have an intrinsic reddening E(B-V) smaller than 3. For each galaxy the minimum allowed galactic extinction was fixed to the Galactic value on the line of sight (Schlafly & Finkbeiner, 2011). The AGN continuum power law index $\alpha$ is the most difficult non stellar parameter to estimate. Indeed, in our very small wavelength domain, variations of $\alpha$ only induce a very weak variation in the slope of the fitted continuum, as explained in Frémaux et al. (2007). We therefore fixed $\alpha$ to the typical value of 1.5 for AGNs222the power law is defined by $\nu^{-\alpha}$ (Vanden Berk et al., 2001). However, these bounds can easily be modified for the synthesis of individual objects.
@@
Such a method is inherently prone to errors due to the presence of noise in the data and potential degeneracies inherent to the studied problem. To assess the impact of these errors on the estimated parameters, we performed a series of Monte Carlo simulations. Gaussian random noise with zero mean and a standard deviation of 10% of the observed galactic spectrum value was added to each pixel of the spectrum. This corresponds approximately to the signal-to-noise ratio of the data. For each object, 300 independent Monte Carlo trials were conducted. To mitigate the influence of outliers, 38% of the highest and 38% of the lowest values for each parameter were excluded. This percentage corresponds to the optimal truncated estimator for the worst-case scenario, that is noise with an undefined mean (Cauchy noise). The standard deviation of the remaining values provides the error bars for the parameters obtained via the minimisation method.
@@
Whatever the approach in stellar syntheses, a spectral database is needed to build a synthetic galactic spectrum. The stellar libraries come in two flavours, either built with theoretical or empirical spectra. Though theoretical libraries could allow for a wide range and combination of stellar parameters, it requires exhaustive model atmospheres that depend on atomic and molecular line opacity data bases not always available.
@@
As demonstrated by Frémaux et al. (2006), the comparison between observed and computed spectra deteriorates for decreasing temperature, where several molecular as well as metallic absorption lines are still missing in atmospheric models. Working in the IR H-band, we thus use an empirical stellar library. It should be noted that each individual spectral type in a library encompasses a significantly large domain of temperature, gravity, and metallicity, because one cannot have as many types of stars in the database as those forming a galaxy.
@@
In the text, we define as young, intermediate and old stellar populations those with stars with ages $\leq$ 107 yrs (supergiant stars), around 108 yrs (A0-F2V), and older than 109 yrs (dwarfs and giants) respectively. The supergiant stars are indeed used as a proxy for a young population since no stars hotter than A type can be considered; the contribution of O and B type stars is negligible in the IR, compared to cold or intermediate stars (A, F spectral types). For more details on the spectral fitting code and database see Frémaux et al. (2007).
Table 2: Stellar parameters.